OpenStudy (anonymous):

Help me please ! Barbara opened a store credit card to purchase a game console for $396. She put the entire purchase on the credit card. Her APR is 24.99% and the minimum payment is 5%. Assuming she makes only minimum payments, what is the remaining balance on her card after 3 months? $362.35 $339.52 $384.45 $351.78

OpenStudy (amistre64):

\[B_n=B_o(1+apr/12)-P\]

OpenStudy (experimentx):

hello \( NaOH + HCl - H_20 \)

OpenStudy (amistre64):

since the payment is 5% of the Balance \[P=B_o(1+apr/12)*.05\] \[B_n=B_o(1+apr/12)-B_o(1+apr/12)*.05\] \[B_n=B_o(1+apr/12)(1-.05)\]

OpenStudy (amistre64):

\[B_1=396(1+\frac{.2499}{12})(.85)\] \[B_2=B_1(1+\frac{.2499}{12})(.85)\] \[B_3=B_2(1+\frac{.2499}{12})(.85)\]

OpenStudy (amistre64):

ugh ... .95 not .85 :)

OpenStudy (anonymous):

how you get (.85) ?

OpenStudy (amistre64):

thats easy ... 1 - .05 = .85 when you dont know how to subtract .... i worked out the recurrsion above in terms of balance and .05 minimum payment

OpenStudy (amistre64):

the new balance part of a bill is: old balance, times (1+apr/12) we subtract a minimum payment of .05 from it old balance, times (1+apr/12); minus old balance, times (1+apr/12), times .05 \[B_n=B_{n-1}(1+r)-B_{n-1}(1+r)(.05)\] \[B_n=B_{n-1}(1+r)[1-.05]\]

OpenStudy (anonymous):

idk how you get .85

OpenStudy (anonymous):

its 0.95

OpenStudy (amistre64):

i know its .95 .... seeing how i am a human being, i tend to miscalculate some simple process. how did i get 1-.05 = .85? its either a typo since 9 and 8 are very close on the keyboard .... or a simply did the subtraction wrong in my head.

OpenStudy (amistre64):

the important thing is: a correction is made and we can move on towards a suitable solution

OpenStudy (anonymous):

k

OpenStudy (amistre64):

\[B_0 = 396\] \[B_1 = 396(.95)(1+\frac{.2499}{12})\] \[B_2 = (396(.95)(1+\frac{.2499}{12}))(.95)(1+\frac{.2499}{12})\\~~~~=396(.95)^2(1+\frac{.2499}{12})^2\] ... \[B_n = 396(.95)^n(1+\frac{.2499}{12})^n\]looks suitable

OpenStudy (amistre64):

the question i have is: do we take the balance of the 3rd or the 4th?

OpenStudy (amistre64):

start: 396 make a payment, remaining balance is 385 make a payment, remaining balance is 373 make a payment, remaining balance is 362 make a payment, remaining balance is 351

OpenStudy (amistre64):

also, is the minimum payment 5% of the original purchase? or is it 5% of the remaining balance?

OpenStudy (amistre64):

its usually a % of the remaining balance ...